# A Unified Spatiotemporal Framework for Modeling Censored and Missing Areal Responses

### v1

@@ 9 Appendix
Let $\boldsymbol{\lambda}=(\rho,\psi,\phi)$ and $T^{*}$, the parameters associated to the covariance structure. The posterior $\pi(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda},T^{*}\mid\mathbf{y})$ can be rewritten as,
$\pi(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda},T^{*}\mid\mathbf{y})\propto\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda},T^{*},\mathbf{y})\pi(\boldsymbol{\lambda},T^{*}\mid\mathbf{y}).$
@@
$\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda},\mathbf{y})\pi(\boldsymbol{\lambda},T^{*}\mid\mathbf{y})\propto\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda})\pi(\boldsymbol{\lambda})\text{L}(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda},T^{*}\mid\mathbf{y}),$
@@
$\pi(\boldsymbol{\lambda},T^{*}\mid\mathbf{y})\propto\frac{\pi(\boldsymbol{\beta},\sigma\mid\boldsymbol{\lambda},T^{*})\text{L}(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda},T^{*}\mid\mathbf{y})}{\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda},\mathbf{y})}\pi(\boldsymbol{\lambda},T^{*}).$ (20)
@@
$\pi(\boldsymbol{\beta},\sigma^{2}|\boldsymbol{\lambda},T^{*},\mathbf{y})$ $\propto$ $\frac{1}{|V_{\boldsymbol{\hat{\beta}}}|^{\frac{1}{2}}}\text{exp}\left[-\frac{\left(\boldsymbol{\beta}-\boldsymbol{\hat{\beta}}\right)^{\top}V^{-1}_{\boldsymbol{\hat{\beta}}}\left(\boldsymbol{\beta}-\boldsymbol{\hat{\beta}}\right)}{2\sigma^{2}}\right]\times$ (21)
@@
$\text{L}(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda},T^{*}\mid\mathbf{y})$ $\propto$ $\frac{1}{\left(\sigma^{2}\right)^{\nicefrac{{n}}{{2}}}\left|\boldsymbol{\Sigma_{\psi}}\right|^{\nicefrac{{1}}{{2}}}}\text{exp}\left[-\frac{(\mathbf{y}-\mathbf{X}\boldsymbol{\beta})^{\top}\boldsymbol{\Sigma_{\psi}}^{-1}(\mathbf{y}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}}\right]$ (22)
$\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda},T^{*})\pi(\boldsymbol{\lambda},T^{*})$ $\propto$ $\frac{\pi(\boldsymbol{\lambda},T^{*})}{\sigma^{2}},$ (23)
@@
$\pi(\boldsymbol{\lambda},T^{*}\mid\mathbf{y})\propto\left|\boldsymbol{\Sigma_{\psi}}\right|^{\nicefrac{{1}}{{2}}}\left|V_{\boldsymbol{\hat{\beta}}}\right|^{\nicefrac{{1}}{{2}}}\left(S^{2}\right)^{-\nicefrac{{n-p}}{{2}}}\pi(\boldsymbol{\lambda,T^{*}})\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\qed$

### v2

@@ 9 Appendix
Let $\boldsymbol{\lambda}=(\rho,\psi,\phi)$ the parameters associated to the covariance structure. The posterior $\pi(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda}\mid\mathbf{y})$ can be rewritten as,
$\pi(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda}\mid\mathbf{y})\propto\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda},\mathbf{y})\pi(\boldsymbol{\lambda}\mid\mathbf{y}).$
@@
$\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda}\mathbf{y})\pi(\boldsymbol{\lambda}\mid\mathbf{y})\propto\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda})\pi(\boldsymbol{\lambda})\text{L}(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda}\mid\mathbf{y}),$
@@
$\pi(\boldsymbol{\lambda}\mid\mathbf{y})\propto\frac{\pi(\boldsymbol{\beta},\sigma\mid\boldsymbol{\lambda})\text{L}(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda}\mid\mathbf{y})}{\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda}\mathbf{y})}\pi(\boldsymbol{\lambda}).$ (20)
@@
$\pi(\boldsymbol{\beta},\sigma^{2}|\boldsymbol{\lambda},\mathbf{y})$ $\propto$ $\frac{1}{|V_{\boldsymbol{\hat{\beta}}}|^{\frac{1}{2}}}\text{exp}\left[-\frac{\left(\boldsymbol{\beta}-\boldsymbol{\hat{\beta}}\right)^{\top}V^{-1}_{\boldsymbol{\hat{\beta}}}\left(\boldsymbol{\beta}-\boldsymbol{\hat{\beta}}\right)}{2\sigma^{2}}\right]\times$ (21)
@@
$\text{L}(\boldsymbol{\beta},\sigma^{2},\boldsymbol{\lambda}\mid\mathbf{y})$ $\propto$ $\frac{1}{\left(\sigma^{2}\right)^{\nicefrac{{n}}{{2}}}\left|\boldsymbol{\Sigma_{\psi}}\right|^{\nicefrac{{1}}{{2}}}}\text{exp}\left[-\frac{(\mathbf{y}-\mathbf{X}\boldsymbol{\beta})^{\top}\boldsymbol{\Sigma_{\psi}}^{-1}(\mathbf{y}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}}\right]$ (22)
$\pi(\boldsymbol{\beta},\sigma^{2}\mid\boldsymbol{\lambda})\pi(\boldsymbol{\lambda})$ $\propto$ $\frac{\pi(\boldsymbol{\lambda})}{\sigma^{2}},$ (23)
@@
$\pi(\boldsymbol{\lambda}\mid\mathbf{y})\propto\left|\boldsymbol{\Sigma_{\psi}}\right|^{\nicefrac{{1}}{{2}}}\left|V_{\boldsymbol{\hat{\beta}}}\right|^{\nicefrac{{1}}{{2}}}\left(S^{2}\right)^{-\nicefrac{{n-p}}{{2}}}\pi(\boldsymbol{\lambda})\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\qed$